Write an equation then solve the inequality


Nine more than half the number n is no more than -8. Find n

Answer by Alan3354(69443)

Write an equation then solve the inequality
-----
An equation is not an inequality.
---
Write an equation
x^2 + 2y^2 - z^2 = 7
That's an equation. It has an equal sign.
================
Nine more than half the number n is no more than -8. Find n
n/2 + 9 <= -8
That's an inequality.
No equal sign, <= instead
----
n/2 + 9 <= -8
Subtract 9
n/2 <= -17
Multiply by 2
n <= -34
=======================
Equations and inequalities are solved in similar manners, but there are differences. Multiplying or dividing by a negative number is one difference.